# Banyan: Improved Representation Learning with Explicit Structure

### v3

@@ 5.4 Ablations
We have shown that Banyan is more effective than its Self-StrAE predecessor, but to tease apart the impact of proposed changes, we perform ablations on our results from our first set of experiments on English (see Table 5).
The simplest impacts stem from the introduction of the diagonalised composition and decomposition functions. These are sigmoided scalar values with which we multiply embeddings and act similarly to the fast weights of (Ba et al., 2016), decaying in the influence of embeddings further down in the tree on the root representation. This means that the embeddings produced by the model are restricted to conform to the compression order dictated by the structure, and we know from (Opper et al., 2023b), that the more we can enforce this constraint the better our representations will end up. Secondly, such simple message passing functions bias the representation space towards informative separability. There has to be some signal from which to perform reconstruction, and all the pressure is now on the representations.
Switching the objective to cross entropy over the vocabulary rather than the contrastive objective used by (Opper et al., 2023b) also yields significant benefits. This is likely because the contrastive loss is supposed to be beneficial because it enforces a pressure for representations to be uniformly distributed in space (Wang & Isola, 2020). However, our other modelling changes already push towards this quality. While it is a shame because the contrastive loss is conceptually elegant. It is known to have problems and eventually lead to shortcut solutions (Robinson et al., 2021). Therefore having a more robust objective grounded in data, like the cross entropy over the vocabulary, is actually quite nice.
@@
Finally, changing to entangled trees is also beneficial. The effect is more pronounced before switching to the cross entropy objective, as it removes the issue of false negatives as discussed in Section 4. However, it also is beneficial beyond this. Entangling explicitly allows the model to take advantage of shared constituency structure between complex sequences, as it combines the information from all incoming parent messages. The fact that performance improves using the cross entropy objective shows that explicitly allowing the model to take advantage of such systematicity is useful.
@@ Appendix A The k and u balance
The change to diagonal composition functions allows us to reduce the number of total parameters while maintaining performance. This is because the number of parameters is directly proportional to channel size u. We show ablations for this finding in Table 7. Our findings are similar to those of (Opper & Siddharth, 2024) the smaller the channel size the better the model performs, although in our case we keep things stable between seeds whereas for them when they simplified they faced issues with extreme instability during training. This is thanks to the new message passing functions.
@@
4 64 42.9 $\pm$ 0.01 43.5 $\pm$ 0.04
8 32 43.2 $\pm$ 0.02 48.6 $\pm$ 0.01
16 16 47.02 $\pm$ 0.03 62.2 $\pm$ 0.01
32 8 49.2 $\pm$ 0.01 62.9 $\pm$ 0.01
64 4 48.7 $\pm$ 0.01 62.9 $\pm$ 0.01
128 2 49.61 $\pm$ 0.02 62.7 $\pm$ 0.001
256 1 48.7 $\pm$ 0.01 62.9 $\pm$ 0.001

### v4

@@ 7.2 Ablations
(a) Uniformity ($\downarrow$) and Alignment loss ($\downarrow$) analysis of Banyan’s representations using: CONTRASTIVE, CE and CE+DIAG.
(b) UMAP analysis of COMP and DECOMP embeddings for a 30k node entangled tree learned by CE w/o diagonal functions.
(c) UMAP analysis of COMP and DECOMP embeddings for a 30k node entangled tree learned by CE with diagonal functions.
Figure 5: Representation Analysis - UMAP Parameters in Appendix C.
@@
Why is Banyan so much more effective than its Self-StrAE predecessor? To probe the impact of each change, we perform ablations using the English STS tasks (Table 6).
@@
Beyond improving efficiency, changing to entangled trees yields some benefits in terms of performance. The effect is significantly more pronounced when using the contrastive objective, as it removes the issue of false negatives as discussed in Section 4. However, it also yields some slight benefit with the CE leaf reconstruction objective. Entangling explicitly allows the model to take advantage of shared constituency structure between complex sequences, as it combines the information from all incoming parent messages. The slight edge this provides indicates that explicitly allowing the model to take advantage of such systematicity may be useful. However, in terms of performance, we find that the choice of functions and objective plays a much bigger role than entangling.
Diagonal Functions: Perhaps the clearest benefits come from the introduction of the diagonalised composition and decomposition functions. These are bounded scalar values (sigmoid) multiplied with embeddings to mimic the time mixing blocks of SSMs (Gu & Dao, 2024). Hierarchically decaying in the influence of embeddings further down the structure through repeated application. This means that the representations are restricted to conform to the compression order it dictates, and we know from (Opper et al., 2023b) that the more we enforce this constraint the better our representations. Secondly, such simple message passing functions bias the representation space towards informative separability. There has to be some signal from which to perform reconstruction, and all the pressure is now on the representations. This is beneficial with the contrastive loss, but really shines when we combine them with cross entropy.
Changing Objective: Our more instructive finding is that cross entropy now outperforms the contrastive loss used by Opper et al. (2023b), contrary to the earlier result that the contrastive objective was critical. Figure 5 provides an analysis with Figure 5(a) showing uniformity vs. alignment metrics ($\downarrow$) from Wang & Isola (2020). These measure (a) how evenly spread embeddings are (uniformity) and (b) the proximity of locally contextual composition embeddings to their globally contextualised counterparts (alignment). Success requires having low scores on both. We can see that the contrastive loss does well, while cross entropy achieves high uniformity, but not alignment without diagonal functions. This is further confirmed when we look at the UMAP (McInnes et al., 2018) analysis of Banyan’s embeddings with (5(c)) and without (5(b)) diagonal functions. Without diagonal functions, composition and decomposition embeddings largely occupy separate subspaces, failing to amortise over context. Introducing diagonal functions results in incredibly high overlap between the two, indicating effective amortisation. We believe that this is due to their simplicity and scaling constraints, restricting complex transformations during message passing. This implicitly forces representations to optimise the same useful qualities as the contrastive loss, without its propensity for shortcut solutions (Robinson et al., 2021). As a result, we can switch objective without compromising representation quality.
@@ 8 Conclusion, Limitations and Future Work
Acknowledgements
MO was funded by a PhD studentship through Huawei-Edinburgh Research Lab Project 10410153. We also wish to thank Henry Conklin, Seraphina Goldfarb-Tarrant, Vivek Iyer, Ivan Vegner, Sahil Verma, Su Kara, Ivan Titov and Edoardo Ponti for their valuable comments and feedback, throughout the research and development of this work. We’d also like to thank the ICML reviewers for their suggestions to help improve the paper. Finally, a special thank you goes to James Morrison, for helping to think through the ideas and providing invaluable feedback and edits to the manuscript.
@@ Appendix A The k and u balance
The change to diagonal composition functions allows us to reduce the number of total parameters while maintaining performance. This is because the number of parameters is directly proportional to channel size u. We show ablations for this finding in Table 7. Our findings are similar to those of (Opper & Siddharth, 2024) the smaller the channel size the better the model performs, although in our case we keep things stable between seeds whereas before simplification caused issues with extreme instability during training. This is thanks to the new message passing functions.
@@
4 64 47.83 $\pm$ 0.2 55.50 $\pm$ 0.22
8 32 47.41 $\pm$ 1.0 62.58 $\pm$ 0.11
16 16 48.01 $\pm$ 1.1 62.73 $\pm$ 0.1
32 8 47.65 $\pm$ 1.1 62.79 $\pm$ 0.07
64 4 48.48 $\pm$ 0.7 62.63 $\pm$ 0.16
128 2 48.53 $\pm$ 1.33 62.97 $\pm$ 0.23
256 1 49.15 $\pm$ 0.6 62.61 $\pm$ 0.23
Appendix B Hyperparameters:
We trained Self-StrAE and Banyan for 15 epochs (circa 15k steps and sufficient for convergence) using the Adam optimiser (Kingma & Ba, 2015), with a learning rate of 1e-3 for Banyan and 1e-4 for Self-StrAE  using a batch size of 512. To process the graphs we used DGL (Wang et al., 2020). The GloVe baseline was trained for 15 epochs with a learning rate of 1e-3, and a window size of 10. We used the official C++ implementation. RoBERTa medium was trained for 200,000 steps, (10% of which were used for warmup). We used a learning rate of 5e-5, and a linear schedule. Positional embeddings are relative key-query. We used the Transformers library to implement and train the model (Wolf et al., 2020). For SimCSE training, we used the default parameters and the official implementation for unsupervised RoBERTa training from Gao et al. (2021). For Sent2Vec we used their official implementation and recommend hyperparameters.
Appendix C UMAP Parameters:
For the UMAP visualisations, we set number of neighbours to 100, minimum distance to 0.3, the metric to cosine and local connectivity to 3. However, the same patterns can we observed through a wide array of hyperparameters, and when changing the metric to euclidean distance. The behavioural changes induced by the diagonal functions remain clear. We selected the above purely based on aesthetic preference for the resulting plots.